# Imperfect quantum networks with tailored resource states

### v1

@@ Appendix A Measurement patterns
In this section we aim to show explicitly the equivalence between strategies introduced in Mor-Ruiz and Dür (2023) and measurement patterns presented in Sec. III. We also present the required tools to translate one to another, and lastly, we analyze the treatment of noisy graph states manipulated via a measurement pattern.
@@ A.1 Preliminaries
Here we present the local Pauli measurements on graph states introduced in Hein et al. (2004, 2006). Consider $v$ to be the qubit of graph state $\left|G\right\rangle$ to be measured in a Pauli basis, then the following unitaries are defined:
@@
$U_{x,+}^{(v)}$ $=\sqrt{i\sigma_{y}^{(w_{0})}}\prod_{w\in\mathcal{N}_{v}-\mathcal{N}_{w_{0}}-\{w_{0}\}}\sigma_{z}^{(w)},$ (9)
$U_{x,-}^{(v)}$ $=\sqrt{-i\sigma_{y}^{(w_{0})}}\prod_{w\in\mathcal{N}_{w_{0}}-\mathcal{N}_{v}-\{v\}}\sigma_{z}^{(w)}.$
@@
The resulting graph $G^{\prime}$ after the measurements is given by the rules described in Sec. II, and after applying the corresponding correction operator Eq. (7), Eq. (8) or Eq. (9), we write the resulting graph state as
@@ A.3 Translation between strategies and measurement patterns
$\sigma_{y}\sqrt{-i\sigma_{z}}$ $=\sqrt{-i\sigma_{z}}\sigma_{x},$
$\sigma_{x}\sqrt{i\sigma_{y}}$ $=-\sqrt{i\sigma_{y}}\sigma_{z},$
$\sigma_{z}\sqrt{i\sigma_{y}}$ $=\sqrt{i\sigma_{y}}\sigma_{x}.$
@@
In Sec. V we use vector $\boldsymbol{m}$ to characterize the measurement pattern of the manipulation of the inner neighbors of a 1D cluster to achieve a Bell pair. This is an instance of the general vector $\bm{\mu}$ presented here. Note that $\bm{\mu}$ and the procedure to calculate it is valid for any graph structure and manipulation.
@@ A.4 Manipulation of noisy graph states with measurement patterns
First, a strategy is formed by a set of manipulation operators that map a graph state to another one, e.g., a local Pauli measurement followed by the corresponding correction operator. When applying a strategy to a noisy graph state, the NSF gives an update rule on how the noise changes after each manipulation operator, so by applying the sequence of manipulation operations to (i) the graph and (ii) the noise operators, step by step, one can retrieve the final graph and the corresponding noise operators. Note that the update rules have been already computed in Mor-Ruiz and Dür (2023) and coded in Wallnöfer (2024) for a set of manipulation operators that are frequently used.
@@
The fact that several strategies might translate to the same measurement pattern is key for the search for the optimal measurement pattern to manipulate an $N$-qubit 1D cluster into a Bell pair between the ends presented in Sec. V.1. All possible measurement patterns are $2^{N-2}$, as one has to count all the possible combinations of $\sigma_{y}$ and $\sigma_{x}$ measurements. Note that, in this case, there is no measurement pattern containing $\sigma_{z}$, since all measurements in a measurement pattern commute, such that a $\sigma_{z}$ measurement could be performed first, leading to a cut in the 1D cluster that would prevent the manipulation to achieve a Bell pair between the ends. In order to perform the exhaustive search, we take an equivalent strategy for each possible measurement pattern and we simulate the corresponding scenario. Importantly, here the use of measurement patterns is again an advantage, because if we searched directly over possible strategies we would have to do so over $3^{N-2}\cdot(N-2)!$ strategies (neglecting the choice of the special neighbor for the $\sigma_{x}$ measurement) and some would not retrieve a Bell pair, but two not entangled qubits instead.
@@ B.1 Local protocol
where $\tau_{n_{i}}$ is the processing time of qubit in $n_{i}$ and $\nu$ is the communication speed, such that $d_{a,n_{i}}/\nu$ is the time for the request started by $a$ to be relayed by all the necessary nodes. As mentioned in Sec. III, in the manipulation there are the inner and outer neighbors. Since each of the two outer neighbors is connected to either $a$ or $b$, we label them as $n_{a}$ and $n_{b}$, and we consider $l$ inner neighbors so that we label them as $\{n_{1},n_{2},\dots,n_{l}\}$. Here, the delay time of the targets includes (i) the time to relay the request to all involved nodes, (ii) the time to perform the measurements and send the outcomes to both $a$ and $b$, and (iii) the time for the other target to finish (i) and (ii). The time to complete (i) and (ii) is
@@ Appendix C Achievable fidelities in a symmetric scenario
Considering the scenario described in Sec. V.1, in Fig. 12, we plot the fidelity of a Bell pair achieved via the manipulation of $N$-qubit 1D cluster, where the inner neighbors are measured using the optimal measurement. Note that depending on the strength of the depolarizing noise and the dephasing time, the chosen measurement pattern changes according to the results presented in Sec. V.1.
(a) Local protocol
(b) Central protocol
Fig. 12: Best achievable fidelity for the local and the central protocol, with even $N$, in terms of the strength of the depolarizing noise $p$ and the dephasing time $T$. The white area denotes that the fidelity is below 0.5, so the target Bell pair is no longer entangled.

### v2

@@ Appendix A Measurement patterns
In this section we aim to show explicitly the equivalence between strategies introduced in [25] and measurement patterns presented in Sec. 3. We also present the required tools to translate one to another, and lastly, we analyze the treatment of noisy graph states manipulated via a measurement pattern.
@@ A.1 Preliminaries
Here we present the local Pauli measurements on graph states introduced in [26, 27]. Consider $v$ to be the qubit of graph state $\left|G\right\rangle$ to be measured in a Pauli basis, then the following unitaries are defined:
@@
$U_{x,+}^{(v)}=\sqrt{i\sigma_{y}^{(w_{0})}}\prod_{w\in\mathcal{N}_{v}-\mathcal{N}_{w_{0}}-\{w_{0}\}}\sigma_{z}^{(w)},\quad\qquad U_{x,-}^{(v)}=\sqrt{-i\sigma_{y}^{(w_{0})}}\prod_{w\in\mathcal{N}_{w_{0}}-\mathcal{N}_{v}-\{v\}}\sigma_{z}^{(w)}.$ (9)
@@
The resulting graph $G^{\prime}$ after the measurements is given by the rules described in Sec. 2, and after applying the corresponding correction operator Eq. (7), Eq. (8) or Eq. (9), we write the resulting graph state as
@@ A.3 Translation between strategies and measurement patterns
$\sigma_{y}\sqrt{-i\sigma_{z}}$ $=+\sqrt{-i\sigma_{z}}\sigma_{x},$
$\sigma_{x}\sqrt{+i\sigma_{y}}$ $=-\sqrt{+i\sigma_{y}}\sigma_{z},$
$\sigma_{z}\sqrt{+i\sigma_{y}}$ $=+\sqrt{+i\sigma_{y}}\sigma_{x}.$
@@
In Sec. 5 we use vector $\bm{m}$ to characterize the measurement pattern of the manipulation of the inner neighbors of a 1D cluster to achieve a Bell pair. This is an instance of the general vector $\bm{\mu}$ presented here. Note that $\bm{\mu}$ and the procedure to calculate it is valid for any graph structure and manipulation.
@@ A.4 Manipulation of noisy graph states with measurement patterns
First, a strategy is formed by a set of manipulation operators that map a graph state to another one, e.g., a local Pauli measurement followed by the corresponding correction operator. When applying a strategy to a noisy graph state, the NSF gives an update rule on how the noise changes after each manipulation operator, so by applying the sequence of manipulation operations to (i) the graph and (ii) the noise operators, step by step, one can retrieve the final graph and the corresponding noise operators. Note that the update rules have been already computed in [25] and coded in [57] for a set of manipulation operators that are frequently used.
@@
The fact that several strategies might translate to the same measurement pattern is key for the search for the optimal measurement pattern to manipulate an $N$-qubit 1D cluster into a Bell pair between the ends presented in Sec. 5.1. All possible measurement patterns are $2^{N-2}$, as one has to count all the possible combinations of $\sigma_{y}$ and $\sigma_{x}$ measurements. Note that, in this case, there is no measurement pattern containing $\sigma_{z}$, since all measurements in a measurement pattern commute, such that a $\sigma_{z}$ measurement could be performed first, leading to a cut in the 1D cluster that would prevent the manipulation to achieve a Bell pair between the ends. In order to perform the exhaustive search, we take an equivalent strategy for each possible measurement pattern and we simulate the corresponding scenario. Importantly, here the use of measurement patterns is again an advantage, because if we searched directly over possible strategies we would have to do so over $3^{N-2}\cdot(N-2)!$ strategies (neglecting the choice of the special neighbor for the $\sigma_{x}$ measurement) and some would not retrieve a Bell pair, but two not entangled qubits instead.
@@ B.1 Local protocol
where $\tau_{n_{i}}$ is the processing time of qubit in $n_{i}$ and $\nu$ is the communication speed, such that $d_{a,n_{i}}/\nu$ is the time for the request started by $a$ to be relayed by all the necessary nodes. As mentioned in Sec. 3, in the manipulation there are the inner and outer neighbors. Since each of the two outer neighbors is connected to either $a$ or $b$, we label them as $n_{a}$ and $n_{b}$, and we consider $l$ inner neighbors so that we label them as $\{n_{1},n_{2},\dots,n_{l}\}$. Here, the delay time of the targets includes (i) the time to relay the request to all involved nodes, (ii) the time to perform the measurements and send the outcomes to both $a$ and $b$, and (iii) the time for the other target to finish (i) and (ii). The time to complete (i) and (ii) is
@@ Appendix C Achievable fidelities in a symmetric scenario
Considering the scenario described in Sec. 5.1, in Fig. 12, we plot the fidelity of a Bell pair achieved via the manipulation of $N$-qubit 1D cluster, where the inner neighbors are measured using the optimal measurement. Note that depending on the strength of the depolarizing noise and the dephasing time, the chosen measurement pattern changes according to the results presented in Sec. 5.1.
[…]
[width=0.5]Fig_12a.pdf
[…]
[width=0.5]Fig_12b.pdf
Figure 12: Best achievable fidelity for the local and the central protocol, with even $N$, in terms of the strength of the depolarizing noise $p$ and the dephasing time $T$. The white area denotes that the fidelity is below 0.5, so the target Bell pair is no longer entangled.